OpenStudy (maheshmeghwal9):

A block of mass 'm' is placed on a smooth wedge of inclination \(theta\). The whole system is accelerated horizontally so that the block doesn't slip on the wedge. What is the magnitude of force exerted by the wedge on the block?

OpenStudy (dls):

mgtan theta? o.O

OpenStudy (maheshmeghwal9):

no; it is \[\frac{mg}{\cos \theta}.\] o.O

OpenStudy (maheshmeghwal9):

but how ?

OpenStudy (dls):

OH :o gotit

OpenStudy (dls):

see

OpenStudy (maheshmeghwal9):

nice:)

OpenStudy (maheshmeghwal9):

i m ready:)

OpenStudy (maheshmeghwal9):

wt happened?

OpenStudy (dls):

wait :(

OpenStudy (maheshmeghwal9):

ok.

OpenStudy (dls):

=>f=ma =>N=mgcos theta + masin theta =>mgcos^2 theta+ mg sin^2 theta(whole upon)/cos theta =>mg/cos theta  did you get anything o.O

OpenStudy (dls):

I SUCK at explaining nvm

OpenStudy (dls):

mg sin theta*

OpenStudy (maheshmeghwal9):

i also did it first time but wt is this in 3rd step mgcos^2 theta+ mg sin^2 theta(whole upon)/cos theta how did u gt this?

OpenStudy (maheshmeghwal9):

however i understood the rest part.

OpenStudy (maheshmeghwal9):

mgcos^2 theta+ mg sin^2 theta(whole upon)/cos theta how did u gt this?

OpenStudy (dls):

wait !

OpenStudy (maheshmeghwal9):

k

OpenStudy (dls):

idk o.o

OpenStudy (maheshmeghwal9):

o.O

OpenStudy (dls):

sorry :|

OpenStudy (maheshmeghwal9):

np:)

OpenStudy (ghazi):

 see you could have done this too by resolving force and considering concept of pseudo force but it is better not to go through that...here just imagine why a block won't sleep...let's say it exerts no weight (that is mg) then there will be no slipping so in the above figure we can see that weight should be equal to m*g= N cos theta \[m*g= N \cos \theta ..... N= \frac{ mg }{ \cos \theta }\] and it is the normal force that is exerted by wedge on block... hope this helps you :)